OpenStudy (anonymous):

HELP WITH COMPOUND INTEREST Determine when, to the nearest year, $4,000 invested at 5% per year, compounded daily, will be worth $10,000.

OpenStudy (tkhunny):

Have you considered teh standard formula? You'll need a logarithm to solve it. \($4,000\cdot \left(1 + \dfrac{0.05}{365}\right)^{n\cdot 365} = $10,000\) The infamous rule of 72 suggests doubling in around 72/5 = 14.4. We're going a little farther than that, si I'm guessing around 17 years? Just from the very vague Rule of 72s and my hunch about the rest. The actual arithmetic shouldn't be TOO FAR from there.